# organization=Université de Rennes, IRISA – UMR 6074,addressline=263 Av. Général Leclerc, city=Rennes, postcode=35000, country=France

### v1

@@ 2 Induced path polytope of chordal graphs
We first present a technical lemma concerning edge contraction in chordal graphs exhibited in [dupontbouillard2024contractions]. After that, we give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated such constraints come in a polynomial number.
@@
If $G=(V,E)$ is chordal, then $G/F$ also is chordal for any set $F\subseteq E$.
Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$, we simply write $C(W)$ when clear from context. We denote by $\mathcal{P}(G)$ the following system:
@@
$x(V)-y(E)$ $=1$ (8)
The next theorem shows that $\mathcal{P}(G)$ describes the induced path polytope of chordal graphs. Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$) and the extremitie of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient one. At most two green edges belong to a same induced path every blue edge is not contained in any induced path containing some other colored edge.
Theorem 2.2.
A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary:
Proof. $\Rightarrow$ We proceed by showing that the pointed cone $\mathcal{C}(G):=$(5)–(7) is generated by the extended incidence vectors of induced paths. The extreme points of (5)–(8) are scalings of the generators of $\mathcal{C}(G)$ satisfying (8).
We now show that the set of induced paths’ extended induced path generates $\mathcal{C}(G)$. To do so, we proceed by induction on a perfect elimination order of $G$. Initialization of the induction is given by the single vertex graph $G$ for which $\mathcal{C}(G)=\mathds{R}_{+}$.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose $\mathcal{C}(H)$ to be generated by $H$’s induced path extended incidence vectors for $H=G\setminus v$. Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $C(H)$.
By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced path $\Psi$ with coefficients $\lambda$. The proof consists of iteratively modifying $\Psi$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Psi$ and $\lambda$; an induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0. Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}=(x^{*},y^{*})_{H}$. Let us fix an edge $uv\in\delta(v)$, we exhibit a conic combination whose $y_{uv}$ component is equal to $y_{uv}^{*}$, as applying the procedure will preserve the value of $y_{e}$ for the other edge $e\in\delta(v)\setminus\{uv\}$, it can be applied until the conic combination is equal to $(x^{*},y^{*})$ on all components but $x_{v}$.
Let us consider the following subsets of $\Psi$:
@@
We refer to Figure 3 for an illustration of the different types of path considered. To each element of $\Psi^{u}$, the addition of $v$ yields another induced path which contains $\{u,v\}$. This implies that unless $y_{uv}^{\prime}$ is already equal to $y_{uv}^{*}$, a conic combination preserving $(x^{\prime},y^{\prime})_{H}$ and strictly increasing over $y_{uv}$ can be computed. Algorithm 2 describes such an operation.
let $P$ be an element of $\Psi^{u}$
$\alpha=\min\{y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P}\}$
$P^{\prime}=P\cup\{v\}$
$(x^{\prime},y^{\prime})=(x^{\prime},y^{\prime})-\alpha\chi^{P}+\alpha\chi^{P^{\prime}}$
$\lambda_{P}=\lambda_{P}-\alpha$
$\lambda_{P^{\prime}}=\lambda_{P^{\prime}}+\alpha$
Algorithm 2 Reduction of $\Psi^{u}\ \forall u\in N(v)$
We can now suppose that $\Psi^{u}$ is empty (otherwise equality of the conic combination’s $y_{uv}$ component and $y_{uv}^{*}$ is already achieved) which implies the following:
$x^{*}_{u}$ $=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})$ (9)
$y_{uv}^{\prime}$ $=\lambda(\Psi_{u}^{v})$ (10)
Claim 2.3.
If $\Psi_{I}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
Proof. Suppose that $\Psi^{u}_{I}$ is empty. Inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ applied to $(x^{*},y^{*})$ yields: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$. Using equality (9) and (10) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}.$
Suppose now that $\Psi_{I}^{u}$ is non-empty.
let $P$ be an element of $\Psi_{I}^{u}$
let $P_{1}$ and $P_{2}$ be the two different maximal subpath of $P$ ending in $u$.
$\alpha=\min(\lambda_{P},\lambda_{u,v}$)
$\lambda_{P}=\lambda_{P}-\alpha$
Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
$j=i+1\mod 2$
$\lambda_{u,v}=\lambda_{u,v}-\alpha$
$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
Algorithm 3 Removal of $\{u,v\}$ from $\Psi_{u}^{v}$
Algorithm (3) is used to ensure that each element of $\Psi_{u}^{v}$ intersects $\delta(u)$ twice, such a path always contains $uv$ and another edge in $\delta(u)$. When Algorithm (3) cannot be applied, either $\Psi_{I}^{u}$ is empty and Claim (2.3) concludes or $\lambda_{u,v}=0$.
Claim 2.4.
If $\Psi_{2}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
Proof. Suppose that $\Psi_{2}^{u}$ is empty. Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$. Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$. Equalities (9) and (10) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and conclude the proof. ∎
Suppose now that $\Psi_{2}^{u}$ is nonempty, let us consider the clique $K\subseteq N(u)\cap N(v)$ of vertices contained in at least one element of $\Psi_{2}^{u}$. We now consider the partition of $\Psi_{I}^{u}$ into $\Psi_{I_{1}}^{u}$ containing the elements intersecting $K$ and $\Psi_{I_{0}}^{u}$ containing the others. Algorithm 4 permits to remove elements from $\Psi_{I_{1}}^{u}$ and adds elements in $\Psi_{u}^{v}$ or $\Psi_{2}^{u}$.
Let $P$ be in $\Psi_{I_{1}}^{u}$
Decompose $P$ into $P_{1}$ and $P_{2}$ the two maximal subpath of $P$ ending in $u$
$\alpha=\min(y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P})$
$\lambda_{P}=\lambda_{P}-\alpha$
Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
$j=i+1\mod 2$
$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
Algorithm 4 Reduction of $\Psi_{I_{1}}^{u}$
Using Algorithm 4 permits obtaining a conic combination of induced path such that $y_{uv}^{\prime}=y_{uv}^{*}$, or no element of $\Psi_{I}^{u}$ intersects $K$ that is $\Psi_{I}^{u}=\Psi_{I_{0}}^{u}$.
Claim 2.5.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ by construction. By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$. ∎
@@
Claim (2.5) now permits to decompose the elements of $\Psi_{I_{0}}^{u}$ for which a neighbor of $u$ is not complete to $K$ into their two maximal subpaths ending in $v$. We then combine one of these subpaths with an element of $\Psi_{2}^{u}$ to obtain other induced paths.
Algorithm 5 gives the procedure. Note that Algorithm 5 may add elements to $\Psi^{u}$ and $\Psi_{I_{1}}^{u}$ which can be empty again using Algorithm 2 and 4. Note that the fact that Algorithm 2 and 4 strictly increase the value of $y_{uv}^{\prime}$ ensures that the sequence of updates on $\lambda$ and $\Psi$ converges.
Let $P_{0}\in\Psi_{I_{0}}^{u}$ and $P_{2}\in\Psi_{2}^{u}$ be such that the $P_{0}^{\prime}\cup P_{2}$ induces a path where $P_{0}^{\prime}$ is a maximal subpath of $P_{0}$ ending in $u$.
$\alpha=\min(\lambda_{P_{2}},\lambda_{P_{0}})$
$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
$\lambda_{P_{0}}=\lambda_{P_{0}}-\alpha$
$\lambda_{P_{2}\cup P_{0}^{\prime}}=\lambda_{P_{2}\cup P_{0}^{\prime}}+\alpha$
$\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}=\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}+\alpha$
Algorithm 5 Replacing the elements of $\Psi_{I_{0}}^{u}$ so that in the remaining ones the neighbors of $u$ are complete to $K$
Now we suppose that the elements of $\Psi^{u}_{I_{0}}$ are complete to $K$ (otherwise, $\Psi^{u}_{2}$ is empty and Claim 2.4 concludes).
Using another time Claim 2.5, we get rid of elements in $\Psi_{u}^{v}$ for which the neighbor of $u$ distinct from $v$ is not complete to $K$ with Algorithm 6. Note that Algorithm 6 increments $\lambda_{uv}$ which is then decreased back using Algorithm 3. It also increments $\Psi_{I_{1}}^{u}$, which is then decreased back using Algorithm 4 if a path in which a neighbor of $u$ is not complete to $K$ is added. Such a sequence of modifications strictly increases $y_{uv}^{\prime}$, ensuring convergence. Note that the neighbor of $u$ in $P_{i}$ as defined in Algorithm 3 is now complete to $K$ as we already used Algorithm 5.
Let $P_{4}$ be in $\Psi_{u}^{v}\setminus\{u,v\}$ and $P_{2}\in\Psi_{2}^{u}$ such that $P_{4}\cup P_{2}\setminus\{v\}$ induces a path
$\alpha=\min(\lambda_{P_{4}},\lambda_{P_{2}})$
$\lambda_{P_{4}}=\lambda_{P_{4}}-\alpha$
$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
$\lambda_{P_{4}\cup P_{2}\setminus\{v\}}=\lambda_{P_{4}\cup P_{2}\setminus\{v\}}+\alpha$
$\lambda_{u,v}=\lambda_{uv}+\alpha$.
Algorithm 6 The neighbors of $u$ in $G[P]$ for $P\in\Psi^{u}$ are complete to $K$
@@
As $(x^{*},y^{*})$ is a point of $\mathcal{C}(G)$, it satisfies the constraint (5) associated with $K$ and $u$ that is:
$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K))+y^{*}(\delta(u,C(K\cup\{u\})))$ (11)
@@
By construction, each element $P_{4}$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K$ implying that $\lambda_{P_{4}}$ contributes one time in $y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]$. By construction, $y^{*}(\delta(u,C(K\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I}^{u})$ . Using equalities (9) and (10) we get:
$2x_{uv}^{*}-2x_{uv}^{*}\geq 2\lambda(\Psi_{2}^{u})+y^{*}_{uv}+2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})-2\lambda(\Psi_{2}^{u})-2\lambda(\Psi_{I}^{u})-2\lambda(\Psi_{u}^{v})$

### v2

@@ 2 Induced path polytope of chordal graphs
We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions]. We use this result to prove a technical graph theory lemma.
After that, we show that the extended incidence vectors of induced paths of a chordal graph form a Hilbert basis. This allows us to give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated constraints come in a polynomial number.
@@
If $G=(V,E)$ is chordal, then $G/F$ is also chordal for any set $F\subseteq E$.
The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis is constructive; therefore, it is necessary to build induced paths with "good properties," and this is what the following lemma stands for.
Claim 2.2.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths. By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$.
To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate. Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system and show in the next theorem that it is generated by the Hilbert basis formed by the induced paths’ extended incidence vectors.
[…]
@@
Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1. At most two green edges belong to a same induced path; every blue edge is not contained in any induced path containing some other colored edge. This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
2.1 Induced paths of chordal graphs and Hilbert basis
In this section, $G$ denotes a chordal graph, $v$ a simplicial vertex, $u$ a neighbor of $v$ and $H$ the graph obtained by removing $uv$ from $G$. We denote by $\Psi$ a family of vertex subsets of $G$ inducing paths, and consider $\lambda$ to be an integer vector indexed by the induced paths of $G$. The pair $(\Psi,\lambda)$ acts as the pair $(\Gamma,\lambda)$ used in the previous section.
We prove that the extended incidence vectors of induced paths of chordal graphs form a Hilbert basis. Although the proof follows the same inductive scheme as in the case of induced trees, it is substantially more technical. In particular, it relies on a refined decomposition of the family $\Psi$.
We first introduce this decomposition of $\Psi$, then define five operations acting on $(\Psi,\lambda)$. For each operation, we establish a corresponding lemma showing that it preserves the property that all elements of $\Psi$ induce paths of $G$, the nonnegativity of $\lambda$ and the value of $\lambda(\Psi)|_{H}:=\sum_{P\in\Psi}\lambda_{P}\xi^{P}|_{H}$. Moreover, we precise its impact on the coefficients $\lambda$. Finally, these ingredients are combined to complete the proof.
Let us consider the following partition of $\Psi$:
@@ 2.1 Induced paths of chordal graphs and Hilbert basis
Figure 3: Types of induced paths considered in Section 2.1
We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions . We denote by $K_{\Psi}$ the elements of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$:
$K_{\Psi}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}$
$K_{\Psi}$ is considered to be changing accordingly to $\Psi$. We further partition $\Psi_{I}^{u}$ into:
  • $\Psi_{K_{\Psi}}^{u}$ containing the elements of $\Psi$ intersecting $K_{\Psi}$
  • $\Psi_{\overline{K}_{\Psi}}^{u}$ containing the others.
We now define some operations to be applied on ($\Psi$$\lambda)$. We say that such an operation is valid if it preserves the integrality and nonnegativity of $\lambda$, if the vertex subsets added to $\Psi$ induce paths, and if it preserves the value of $\lambda(\Psi)|_{H}$. In each reduction, $\alpha$ is chosen as an integer satisfying $0<\alpha\leq\min\{\lambda_{P}:P\in\Psi,\ \lambda_{P}\text{ is decreased}\}$. This ensures that $\lambda$ stays nonnegative.
$\mathbf{\Psi^{u}}$-Reduction : let $P$ be an element of $\Psi^{u}$. Increase $\lambda_{P\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
$\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{K_{\Psi}}^{u}$. $P$ contains the two inclusion-wise maximal distinct subsets $P_{1}$ and $P_{2}$ inducing paths ending in $u$. Let $P_{2}$ be the one intersecting $N(v)\setminus u$. Increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
$\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$ such that a neighbor $u_{1}$ of $u$ in $P$ is not complete to $K_{\Psi}$. Let $P_{1}$ be its inclusion-wise maximal subset inducing a path ending in $u$ and containing $u_{1}$, and let $P^{\prime}$ be the other inclusion-wise maximal subset inducing a path ending in $u$. Let $P_{2}$ be an element of $\Psi_{2}^{u}$ whose element $u_{2}$ in $K_{\Psi}$ is non-adjacent to $u_{1}$. Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ by $\alpha$.
$\mathbf{\Psi_{u}^{v}}$-Reduction : Let $P$ be a path of $\Psi_{u}^{v}$ such that the neighbor $u_{1}$ of $u$ distinct from $v$ is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$. Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$. Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$.
$\mathbf{\lambda_{\{u,v\}}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$. $P$ contains the two distinct inclusion-wise maximal subset $P_{1}$ and $P_{2}$ both inducing paths ending in $u$. Note that $P_{1}\cup\{v\}$ induces a path. Decrease $\lambda_{u,v}$ and increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$.
Note that in each of the above operations, $\alpha$ is integer, implying that they preserve integrality of $\lambda$. We show that under some conditions, the above reductions are valid. It remains to verify that the sets of vertices built induce paths and that the value of $\lambda(\Psi)|_{H}$ is preserved.
Lemma 2.3.
Suppose that $\Psi^{u}$ is nonempty. $\mathbf{\Psi^{u}}$-Reduction is valid.
Proof. Given $\in\Psi^{u}, PP$ ends in $u$ and does not intersect $N(v)\setminus\{u\}$, therefore $P\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P\cup\{v\}}|_{H}$. ∎
Lemma 2.4.
Suppose that $\Psi_{K_{\Psi}}^{u}$ is non-empty. $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and strictly increments $\lambda(\Psi^{u}_{2})$, $\lambda(\Psi^{v}_{u})$.
Proof. By construction, $P\in\Psi_{K_{\Psi}}^{u}$ is decomposed into its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ that induce paths ending in $u$. At most one of them intersects $N(v)\setminus\{u\}$, which implies that $P_{1}\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}.$ Therefore, $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and increments both $\lambda_{P_{1}\cup\{v\}}$ and $\lambda_{P_{2}}$. ∎
Lemma 2.5.
Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ contains an element $P$ such that a neighbor $u_{1}\in P$ of $u$ is not complete to $K_{\Psi}$, in particular $\Psi_{2}^{u}$ is non empty. Then $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
Proof. Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$, let $P^{\prime}$ be its other subset inducing a path ending in $u$. Let $P_{2}$ be an element of $\Psi^{u}_{2}$ such that the neighbor $u_{2}$ of $u$ in $P_{2}$ is non adjacent to $u_{1}$. Claim 2.2 implies that $P_{1}\cup P_{2}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}\cup P_{2}}|_{H}+\xi^{P^{\prime}\cup\{v\}}|_{H}.$ Therefore $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid. ∎
Lemma 2.6.
Suppose that $\Psi_{u}^{v}$ contains an element $P$ such that the neighbor $u_{1}$ of $u$, distinct from $v$, is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$. $\mathbf{\Psi_{u}^{v}}$-Reduction is valid and strictly increases $\lambda_{\{u,v\}}$.
Proof. Let $P_{1}=P\setminus\{v\}$, by Claim 2.2 $P_{1}\cup P_{2}$ induces a path and $\{u,v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}}|_{H}+\xi^{\{u,v\}}|_{H}$. ∎
Lemma 2.7.
Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non empty. $\mathbf{\lambda_{\{u,v\}}}$-Reduction is valid and strictly decreases $\lambda_{\{u,v\}}$.
Proof. Given $P\in\Psi_{\overline{K}_{\Psi}}^{u}$, we consider its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ inducing a path ending in $u$. $P_{1}\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{\{u,v\}}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}$. ∎
The proof of the following theorem proceeds by induction over the number of vertices and makes use of the previous definitions and operations.
Theorem 2.8.
The extended incidence vectors of induced paths of a chordal graph form a Hilbert basis of $\mathcal{C}_{P}(G)$.
Proof. The statement is trivial for $|V|=1$. Induction goes as follows: As $G$ is chordal, it contains a simplicial vertex $v$. The graph $G\setminus v$ is chordal and by hypothesis, the induced paths’ extended incidence vectors of $G\setminus v$ form a Hilbert basis of $\mathcal{C}_{P}(G\setminus v)$. Let $(x^{*},y^{*})$ belong to $\mathcal{C}_{P}(G)\cap\mathds{Z}^{V\cup E}$, its restriction $(x^{*},y^{*})|_{G\setminus v}$ belongs to $\mathcal{C}_{P}(G\setminus v)\cap\mathds{Z}^{V\cup E}$. From now on, $(\Psi,\lambda)$ denotes an arbitrary family of induced paths of $G$ with integer coefficients satisfying $(x’,y’)|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$, where (x,y)(x’,y’) is the point induced by $(\Psi,\lambda)$. All operations defined above transform $(\Psi,\lambda)$ into another such family and preserve: integrality of $\lambda$, the fact that the elements of $\Psi$ induce paths of $G$ and $(x’,y’)|_{H}=(x^{*},y^{*})|_{H}$. Note that the latter requirement implies that the operations not only preserve the value of $\lambda(\Psi)$ over elements in $G\setminus v$ but also on edges in $\delta(v)\setminus\{uv\}$.
The goal of the upcoming sequence of reductions is to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$. For this we consider the slackness;
$\Delta:=\sum_{uv\in\delta(v)}\Delta(u,v)\text{ where }\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}.$
The proof proceeds by fixing a neighbor $u$ of $v$ and modifying $(\Psi,\lambda)$ until the equality $\Delta(u,v)=0$ holds. Since reductions for a fixed $u$ do not affect $\Psi_{w}^{v}$ for $w\neq u$, and that elements of $\Psi^{u}_{w}$ are the only ones contributing to the value of $y_{vw}^{\prime}$, the procedure can be applied independently for each neighbor of $v$.
An induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0.
Lemmas 2.32.5 give conditions for some valid reductions strictly decreasing $\Delta(u,v)$ to be applied for respectively $\alpha=\min(\Delta(u,v),\lambda_{P})$, $\alpha=\min(\Delta(u,v),\lambda_{P})$, and $\alpha=\min(\lambda_{P},\lambda_{P_{2}},\Delta(u,v))$. Since the corresponding reductions strictly decrease either $\Delta(u,v)$, at some point, none of Lemma 2.32.6 apply for the fixed neighbor $u$. This implies that, either $\alpha=0$ and $\Delta(u,v)=0$ which ends the proof or the following holds:
  • $\Psi^{u}$, $\Psi^{u}_{I_{1}}$ both are empty
  • the neighbors of $u$ in the elements of $\Psi^{u}_{I_{0}}$ are complete to $K_{\Psi}$.
By construction, the only elements of $\Psi$ which contribute to the value of $y_{uv}^{\prime}$ belong to $\Psi_{u}^{v}$, therefore the following holds :
$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
As $\Psi^{u}$ is empty, the following holds:
$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{\overline{K}_{\Psi}}^{u})+\lambda(\Psi_{u}^{v})\\$ (9)
Suppose that $\Psi^{u}_{I_{0}}$ is empty. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$. Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta(u,v)=0$.
Assume henceforth that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non-empty.
Suppose $\Psi_{2}^{u}$ to be empty, Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$. Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$. Equalities (9) and (8) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and $\Delta(u,v)=0$.
We may now assume that $\Psi_{2}^{u}$ is nonempty.
Therefore, $K$ is non-empty and the neighbors of $u$ in the elements of $\Psi_{\overline{K}_{\Psi}}^{u}$ are complete to $K_{\Psi}$.
Suppose that the conditions of Lemma 2.6 hold, then applying $\Psi_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$ yields an element of $\Psi_{K_{\Psi}}^{u}$, therefore, the conditions of Lemma 2.4 hold and $\Psi_{K_{\Psi}}^{u}$-Reduction applies. This sequence of operations strictly decreases $\Delta(u,v)$.
Suppose that the conditions of Lemma 2.7 hold, applying $\lambda_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{u,v}$) yields an element of $\Psi_{u}^{v}$, and one of $\Psi^{u}$. Therefore, the conditions of Lemma 2.3 hold, and $\Psi^{u}$-Reduction applies. This sequence of operations also decreases $\Delta(u,v)$.
As $\Psi^{u}$-Reduction, $\Psi_{K_{\Psi}}^{u}$-Reduction, $\Psi_{\overline{K}_{\Psi}}^{u}$-Reduction, and the two latter sequences of reductions strictly decrease $\Delta(u,v)$, we may assume that none of Lemmas 2.32.7 apply or that $\Delta(u,v)=0$. Therefore, the neighbors of $u$ in the elements of $\Psi^{v}_{u}$ are complete to $K_{\Psi}$ and $\lambda_{\{u,v\}}=0$.
@@
As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is (where $C(S)$ denotes the set of vertices complete to $S$):
$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K_{\Psi}))+y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))$ (10)
@@
By construction, each element $P$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K_{\Psi}$. This implies that $\lambda_{P}$ contributes two times in $y^{*}[\delta(u,C(K_{\Psi}\cup\{u\}))]$. By construction, $y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K_{\Psi}\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{\overline{K}_{\Psi}}^{u})$. Using equalities (8) and (9) we get:
$0\geq y^{*}_{uv}-\lambda(\Psi_{u}^{v})$